# Vipera: Blending Visual and LLM-Driven Guidance for Systematic Auditing of Text-to-Image Generative AI

### v1

@@ 3.1.2. Findings
Typical scales and motivations for T2I auditing. All participants reported that the scale of auditing varies based on the specific task. While the most common average scale for auditing was reviewing around 10 images at once, some participants indicated that they had audited hundreds of images simultaneously in certain cases. Motivations for large-scale auditing differed among participants: auditors who were also developers often engaged in systematic testing. For example, P1 reported a pattern of category-based image generation and auditing: “The occupations I used were teacher, dishwasher, janitor, and then there were higher-class occupations like CEO, lawyer, and doctor. I asked the model to generate a series of images for each occupation, and then I used a visual assistant AI to gather demographic data about the generated images. This allowed me to collect information on what the model perceives a janitor and a CEO to look like. I ended up generating around 30 images for each of the 10 occupations, resulting in a total of about 300 images.” In contrast, non-developer auditors preferred smaller scales typical of their daily work but would analyze a larger image corpus if the initial images have issues or did not meet their requirements. As P3 noted, “I often focus on the image quality first… My goals would be generating [ideal] images rather than testing the model.
@@ 3.2. Design goals
DG1: Leverage intuitive and interactive visual aids for structured, customizable auditing and effective result sensemaking. According to the formative study, the visual aids, including the scene graph and the bar charts, might bring cognitive load to users, despite their effectiveness in inspiring audit directions and helping users understand the images. According to the user feedback, we infer two crucial reasons that account for this. First, the graphical layout and the size of the scene graph might overwhelm users, making them confused about where to focus on during the analysis. The other was the lack of interactiveness in the visual aids: the users were forced to reactively check the suggested criteria instead of proactively creating them, limiting their creativity and activeness. Moreover, it is hard for users to associate the statistics back with the images due to the lack of coordinated views. Therefore, Vipera should grant users full agency through diverse interactions while allowing them to navigate between visual cues and data at any time.
DG2: Enable intuitive comparison of how the distribution of large-scale image outputs changes across prompts. Comparing different images has been demonstrated as a rich source of inspiration for auditing in prior literature (Shen et al., 2021; Deng et al., 2025a). Likewise, we hypothesize that comparing the distributions and prompts may similarly provide abundant inspiration for users. To support this, Vipera should facilitate easy comparison of prompts and image distributions while highlighting the differences through proper visual assistance.
@@
DG5: Assist auditors in documenting, synthesizing, communicating, and acting upon their audit findings. The ultimate goal of an audit is to produce a coherent, evidence-backed report that can inform stakeholders and drive action. The process of documenting insights while simultaneously conducting an exploratory analysis is cognitively demanding and often disjointed. To address this, Vipera should integrate documentation directly into the analytical interface. This involves providing tools for auditors to seamlessly capture their thoughts, bookmark key visual evidence (such as charts or specific images), and structure their findings as they emerge, ultimately streamlining the transition from exploration to a communicable, actionable report.
@@ 4. The Vipera System
This section presents the design and implementation details of Vipera. As shown in  Figure 1, three major components are included: a input view for creating prompts and generating images (A), an analysis view (B) for customizing the auditing process, seeking guidance, and analyzing the results, and a note view (C) for report generation.
@@ 4.1.1. Input view
The input view incorporates a input box for writing prompts, a numerical input box for specifying the number of images to be generated, and a model selector for choosing the model to be audited. The user can revisit this view at any time during the auditing to create new prompts or images.
@@ 4.1.2. Analysis view
Images. The image view shows the thumbnails of the generated images, with the border color indicating the corresponding prompt. Users can click on the images to see the full-sized version, where a bookmark button is provided for saving images of interest for later report generation. In addition, they can hover on the image to see the labels for the auditing criteria defined in the scene graph, and meanwhile the associated nodes in the graph will be highlighted in the scene graph (Figure 3). When they notice inaccurate labels, they are also allowed to edit the labels by right-clicking on the image, selecting the edit option in the consequent menu to enter a modal, and manually editing the labels in it.
Scene graph. The scene graph serves as both a visual summary of the semantic contents within the images and a visual cue for users to organize their auditing (DG1). Objects in the scene graph are represented as nodes and organized in a tree layout based on their semantic relationship. The tree layout is used to replace the graphical one in ViperaBase to for better intuitiveness (DG1), helping users organize their thoughts and navigate in the auditing space.
@@
  • Scope type: Since Vipera supports iterative creation of prompts and auditing criteria, the scope of nodes in the scene graph can be dynamic to accmmodate the prompts and images to come. Given this, Vipera supports two scope types: fixed and auto-extended. If an auto-extended scope is specified, the criteria will be applied to newly generated images, which is not the case for a fixed scope.
  • Candidate values (optional): To prevent the LLM from generating overly diversified labels for the images that hardly leads to meaningful distinctions, Vipera allows user to specify a list of predefined candidate label values for classification (e.g., male and female for the attribute node gender). If provided, the LLM will be constrained to choose from this list when labeling images.
When an attribute node is added, the node will be used as a criterium for a labeler LLM to label the images. The result labels will be visualized in stacked bar charts with colors indicating the corresponding prompts to facilitate comparison (DG2). Users can hover on the bars to navigate back to the matched images highlighted in the image view (Figure 3). Additionally, they can edit the information of a node or choose to relabel the images based on an attribute node via the context menu. A relabeling will also be triggered when the user modifies the candidate values or the scope. In this case, two options are provided: relabeling all images or only relabeling images that are affected (i.e., contain a label that is not within the candidate values or not in the scope). The user can change the option in the relabeling modal.
Suggestions. Informed by DG3, Vipera provides two classes of auditing guidance, audit analysis support and prompt suggestion, to inspire users of new auditing criteria and prompts respectively.
@@
  • Prompt suggestion. This module promotes divergence thinking by suggesting insightful prompts, replacing words or phrases in existing prompts. Upon users deciding to try out the prompt, Vipera will duplicate the existing auditing criteria while generating new images, allowing for an effective comparison of results between different prompts. For instance, when the user applies the suggestion replacing the word “doctor” with “nurse” in the prompt “a cinematic photo of a doctor”, Vipera will automatically add an object node “nurse” to the sibling of “doctor” while duplicating the descendent nodes of “doctor” in this new branch. Such a transformation makes sure that the labeling results of newly generated images for nurses will be displayed in the correct branch.
Figure 3. View coordinations between the image view and the scene graph. When the user hovers on an image, a pop-up will appear showing its labels, and the relevant attribute nodes in the scene graph will be highlighted. When the user hovers on a bar in an embedded bar chart in the attribute node, the corresponding images will be highlighted in the image view.:
Figure 4. Vipera’s technical pipeline. Black edges indicate the data flow: (A) The prompt from the user is fed to T2I model to generate images. (B) A scene graph is generated from the images (see Figure 5 for details). (C) Upon an attribute node is added to the scene graph, the images will be labeled using the specified criteria. (D) The labels will be aggregated and visualized as a stacked bar chart. (E) The charts will be embedded into the attribute nodes of the scene graph for rendering. (F) The images and charts will be added to the auditing report upon being bookmarked by the user, with the user’s comments attached to them. Blue edges indicate the inspiration flow for iterative auditing: Prompts and auditing criteria can be inspired by inspecting the images (a), thinking with the scene graph (b), and applying the LLM-driven guidance (c).:
@@ 4.1.3. Note view
To facilitate the authoring of auditing reports (DG5), a note view is incorporated, which consists of two tabs: a tab of General Notes that includes an input box for noting down the overall insights, and a tab of Evidence that shows the bookmarked charts and images. Detailed user comments can be appended to specific charts or images as fine-grained insights. Moreover, Vipera features LLM-powered auto-completion when users write in this view to further enhance the auditing efficiency. The prompts, bookmarked items, and existing notes will be injected into a LLM to generate the completion, which can be easily applied upon users tap on the tab key on the keyboard.
@@ 4.2. User workflow and technical pipeline
Vipera has been developed and deployed as a web-based application. Figure 4 shows the workflow of Vipera [^3] All system prompts used in the pipeline are provided in the appendix.. After generating images from a user’s prompt, Vipera initiates the scene graph construction process, which is detailed in Figure 5. It starts by randomly sampling four images (or all, if fewer than four are available) to serve as a representative basis for the initial scene graph (Figure 5A). While this small sample may not capture the entire corpus, it provides a manageable starting point for the user. For each sampled image, Vipera prompts an omni-modal LLM (Gemini 2.5 Flash in our implementation) to extract a tree-based scene graph that captures the image’s semantics (Figure 5B). The LLM is instructed to identify key physical objects and notable features and organize them into a tree structure. To provide a consistent high-level organization, we constrain the first level of the tree to “foreground” and “background” nodes in our study, which is an intuitive and widely applicable distinction. The resulting scene graphs from the four images are then merged, with their leaf nodes combined (Figure 5C). To avoid overwhelming the user, a maximum of five leaf nodes are randomly selected for the final aggregated scene graph (Figure 5D, DG1).
@@
With the scene graph serving as a visual summary, users can add object or attribute nodes, where attribute nodes will serve as auditing criteria. When an attribute node is added, Vipera begins to evaluate and label the images using this criterium (Figure 4C). First, it updates the scene graph with the new node and ensures that all images within the new node’s scope are also included in the scope of its ancestor nodes up to the root. Then, it extracts a partial graph schema, which is a subset of the scene graph containing the new node and its path to the root. This schema is used to prompt a omni-modal LLM (GPT-5-mini in our implementation) to generate a label for each image within the specified scope. If the user has defined candidate values for the attribute, the LLM is constrained to choose from that list. The resulting labels are then aggregated, and their distribution is visualized as a stacked bar chart (Figure 4D) embedded into the added node using the D3.js library [^4] https://d3js.org/ (Figure 4E). Furthermore, users can bookmark the images and the charts within scene graph nodes at any time during auditing and comment about them in the note view for an auditing report (Figure 4F).
@@
Figure 5. The detailed technical pipeline for generating the initial scene graph. (A) A random subset is sampled from the images. (B) For each image, an individual scene graph is generated through an omni-modal LLM. (C) The result scene graphs are merged into one. (D) The result is pruned by keeping a maximum of five leaf nodes.:
@@
Additionally, users can benefit from diverse sources of inspirations for iterative exploration of prompts and auditing criteria in Vipera, illustrated as blue arrows in Figure 4. Specifically, they may inspect the images carefully to get insights and new inspirations for prompts and criteria (Figure 4a). Such inspirations may also come from the scene graph, where the user either thinks with its structure about potential nodes to be added or examine the embedded charts for insights (Figure 4b). Finally, they may adopt the LLM-powered suggestions at any time for new auditing ideas (Figure 4c). Such an inspiration loop is built upon the artifacts from the data loop and serves as the engine for the user workflow, promoting iterative, mixed-initative, and comprehensive auditing.
@@ 4.3. Usage Scenario
To begin his audit, Bob enters the prompt “A cinematic photo of a doctor” into the input view and generates 15 images (Figure 1, step 1). Vipera populates the analysis view with the prompt and the images, while showing a corresponding scene graph that characterizes the semantics of images (Figure 1, step 2). The graph organizes the content into foreground (e.g., doctor, nurse) and background (e.g., office, medical equipment) nodes.
@@
Afterwards, Bob felt tired to come up with new auditing ideas and decided to refer to the AI suggestions. He first examined the Audit Analysis Support, which highlighted two images with a key difference and suggested adding a stethoscope attribute to the doctor node. Bob found the suggestion insightful and applied the suggestion (Figure 1, step 5). The resulting chart showed that some doctors were not wearing a stethoscope. To verify this, he hovered over the corresponding bar, which highlighted the specific images in the image view (Figure 3), allowing for rapid, coordinated sensemaking.
Next, Bob noticed the Prompt Suggestion recommending he replace “doctor” with “nurse” to explore a related profession. He accepted the prompt, and Vipera generated a new set of images (Figure 1, step 6). Crucially, the system also added a new nurse branch to the scene graph and automatically duplicated the existing criteria (gender and stethoscope) under it. With Vipera’s color-coded bar charts, Bob compared the results of these two prompts on the scene graph easily and acquired new insights in occupational stereotypes. From this point, Bob can continue his audit through multiple iterations of structured exploration and comparison.

### v2

@@ 3.1.2. Findings
Typical scales and motivations for T2I auditing. All participants reported that the scale of auditing varies based on the specific task. While the most common average scale for auditing was reviewing around 10 images at once, some participants indicated that they had audited hundreds of images simultaneously in certain cases. Motivations for large-scale auditing differed among participants: auditors who were also developers often engaged in systematic testing. For example, P1 reported a pattern of category-based image generation and auditing: “The occupations I used were teacher, dishwasher, janitor, and then there were higher-class occupations like CEO, lawyer, and doctor. I asked the model to generate a series of images for each occupation, and then I used a visual assistant AI to gather demographic data about the generated images. This allowed me to collect information on what the model perceives a janitor and a CEO to look like. I ended up generating around 30 images for each of the 10 occupations, resulting in a total of about 300 images.” In contrast, non-developer auditors preferred smaller scales typical of their daily work but would analyze a larger image corpus if the initial images had issues or did not meet their requirements. As P3 noted, “I often focus on the image quality first… My goals would be generating [ideal] images rather than testing the model.
@@ 3.2. Design goals
DG1: Leverage intuitive and interactive visual aids for structured, customizable auditing and effective result sensemaking. According to the formative study, the visual aids, including the scene graph and the bar charts, might bring cognitive load to users, despite their effectiveness in inspiring audit directions and helping users understand the images. According to the user feedback, we infer two crucial reasons that account for this. First, the graphical layout and size of the scene graph may overwhelm users, causing confusion about where to focus during the analysis. The other was the lack of interactiveness in the visual aids: the users were forced to reactively check the suggested criteria instead of proactively creating them, limiting their creativity and activeness. Moreover, it is hard for users to associate the statistics back with the images due to the lack of coordinated views. Therefore, Vipera should grant users full agency through diverse interactions while allowing them to navigate between visual cues and data at any time.
DG2: Enable intuitive comparison of how the distribution of large-scale image outputs changes across prompts. Comparing different images has been demonstrated as a rich source of inspiration for auditing in prior literature (Shen et al., 2021; Deng et al., 2025b). Likewise, we hypothesize that comparing the distributions and prompts may similarly provide abundant inspiration for users. To support this, Vipera should facilitate easy comparison of prompts and image distributions while highlighting the differences through proper visual assistance.
@@
DG5: Assist individual auditors in documenting, synthesizing, communicating, and acting upon their audit findings. The ultimate goal of an audit is to produce a coherent, evidence-backed report that can inform stakeholders and drive action. The process of documenting insights while simultaneously conducting an exploratory analysis is cognitively demanding and often disjointed. To address this, Vipera should integrate documentation directly into the analytical interface. This involves providing tools for auditors to seamlessly capture their thoughts, bookmark key visual evidence (such as charts or specific images), and structure their findings as they emerge, ultimately streamlining the transition from exploration to a communicable, actionable report.
@@ 4. The Vipera System
This section presents the design and implementation details of Vipera. As shown in  Figure 2, three major components are included: an input view for creating prompts and generating images (A), an analysis view (B) for customizing the auditing process, seeking guidance, and analyzing the results, and a note view (C) for report generation.
Figure 2. The Vipera interface. (A) The input box for creating prompts and specifying the number of images. (B) The analysis view for interactive auditing. Left: The prompts and AI-powered suggestions of auditing criteria and prompts. Right: The generated images and an interactive scene graph summarizing the image semantics to inspire users and guide the auditing process. (C) The note view for composing an auditing report.
Figure 1: The interface displays a tool for auditing generative text-to-image models. On the top (A), there is an input view which includes an input box for prompts, a numerical input box for the number of images, a selector for the model to be audited, and a ”Generate” button for starting the image generation. Right below (A) is the analysis view (B). On the left side, a text box shows the prompt ”A cinematic photo of a doctor,” with options for audit analysis support with optional keywords for targeted suggestion generation and prompt suggestions. To the right, the images generated are placed here, annotated with different colors corresponding to the prompts. There is also a scene graph outlines the relationships between elements, including nodes such as ”doctor”, ”nurse”, ”office”, ”medical equipment, and so on.” Finally, there is a note view at the bottom, including a ”General Notes” tab for overall insights and a ”Evidence” tab for bookmarked charts and images. For instance, the current notes in the Evidence tab include a bar chart about doctors’ genders and a comment pointing out the gender bias.
@@ 4.1.1. Input view
The input view incorporates an input box for writing prompts, a numerical input box for specifying the number of images to be generated, and a model selector for choosing the model to be audited. The user can revisit this view at any time during the auditing to create new prompts or images.
@@ 4.1.2. Analysis view
Images. The image view shows the thumbnails of the generated images, with the border color indicating the corresponding prompt. Users can click on the images to see the full-sized version, where a bookmark button is provided for saving images of interest for later report generation. In addition, they can hover on the image to see the labels for the auditing criteria defined in the scene graph, and meanwhile, the associated nodes in the graph will be highlighted in the scene graph (Figure 3). When they notice inaccurate labels, they are also allowed to edit the labels by right-clicking on the image, selecting the ”Edit” option in the subsequent menu to enter a modal, and manually editing the labels within it.
Scene graph. The scene graph serves as both a visual summary of the semantic contents within the images and a visual cue for users to organize their auditing (DG1). Objects in the scene graph are represented as nodes and organized in a tree layout based on their semantic relationship. The tree layout is used to replace the graphical one in ViperaBase for better intuitiveness (DG1), helping users organize their thoughts and navigate in the auditing space.
@@
  • Scope type: Since Vipera supports iterative creation of prompts and auditing criteria, the scope of nodes in the scene graph can be dynamic to accommodate the prompts and images to come. Given this, Vipera supports two scope types: fixed and auto-extended. If an auto-extended scope is specified, the criteria will be applied to newly generated images, which is not the case for a fixed scope.
  • Candidate values (optional): To prevent the LLM from generating overly diversified labels for the images that hardly lead to meaningful distinctions, Vipera allows users to specify a list of predefined candidate label values for classification (e.g., male and female for the attribute node gender). If provided, the LLM will be constrained to choose from this list when labeling images.
Figure 3. View coordinations between the image view and the scene graph. When the user hovers over an image, a pop-up will appear showing its labels, and the relevant attribute nodes in the scene graph will be highlighted. When the user hovers on a bar in an embedded bar chart in the attribute node, the corresponding images will be highlighted in the image view.
The above half of the image shows several images generated by the T2I model, while the bottom half shows the scene graph. The user hovers on an image, and a pop-up appears, showing the labels ”stethoscope: wearing” and ”gender: male”, with a bar chart showing that 87% and 100% percent of images share the same labels in these criteria, respectively. Meanwhile, the nodes ”stethoscope” and ”gender” in the scene graph are highlighted. When the user hovers on the bar of ”not wearing” of the bar chart embedded into the node ”stethoscope”, the images corresponding to this bar are highlighted in the image view.
When an attribute node is added, the node will be used as a criterion for a labeler LLM to label the images. The result labels will be visualized in stacked bar charts with colors indicating the corresponding prompts to facilitate comparison (DG2). Users can hover on the bars to navigate back to the matched images highlighted in the image view (Figure 3). Additionally, in case the labeler itself makes mistakes, users can edit the information of a node or choose to relabel the images based on an attribute node via the context menu. A relabeling will also be triggered when the user modifies the candidate values or the scope. In this case, two options are provided: relabeling all images or only relabeling images that are affected (i.e., contain a label that is not within the candidate values or not in the scope). The user can change the option in the relabeling modal.
Suggestions. Informed by DG3, Vipera provides two classes of auditing guidance, audit analysis support and prompt suggestion, to inspire users of new auditing criteria and prompts, respectively.
@@
  • Prompt suggestion. This module promotes divergent thinking by suggesting insightful prompts, replacing words or phrases in existing prompts. Upon users deciding to try out the prompt, Vipera will duplicate the existing auditing criteria while generating new images, allowing for an effective comparison of results between different prompts. For instance, when the user applies the suggestion replacing the word “doctor” with “nurse” in the prompt “a cinematic photo of a doctor”, Vipera will automatically add an object node “nurse” to the sibling of “doctor” while duplicating the descendent nodes of “doctor” in this new branch. Such a transformation makes sure that the labeling results of newly generated images for nurses will be displayed in the correct branch.
Figure 4. Vipera’s technical pipeline. Black edges indicate the data flow: (A) The prompt from the user is fed to the T2I model to generate images. (B) A scene graph is generated from the images (see Figure 5 for details). (C) Upon adding an attribute node to the scene graph, the images will be labeled according to the specified criteria. (D) The labels will be aggregated and visualized as a stacked bar chart. (E) The charts will be embedded into the attribute nodes of the scene graph for rendering. (F) The images and charts will be added to the auditing report upon being bookmarked by the user, with the user’s comments attached to them. Blue edges indicate the inspiration flow for iterative auditing: Prompts and auditing criteria can be inspired by inspecting the images (a), thinking with the scene graph (b), and applying the LLM-driven guidance (c).
The flowchart illustrates a process for generating and analyzing images using the T2I model. It begins with a prompt that leads to the creation of images. These images are then sampled and extracted to form a scene graph. The scene graph is labeled to produce labeled images, which can be aggregated and visualized in bar charts that will later be embedded into the scene graph. Additionally, there is a section for prompt ideas and comparisons, generated by the LLMs using the prompts, images, and scene graph. Meanwhile, there is also an inspiration flow complementary to the above data flow, where users can get auditing inspirations on prompts and criteria from inspecting images, thinking with the scene graph (or checking the charts embedded), or applying the LLM-powered suggestions, and consequently create more prompts and criteria for iterative auditing. The entire process is driven by an inspiration flow (in blue arrows), which flows from right to left in the diagram. Another data flow (in black arrows, flowing from left to right) is incorporated to generate content that powers the inspiration flow.
@@ 4.1.3. Note view
To facilitate the authoring of auditing reports for individual auditors (DG5), a note view is incorporated, which consists of two tabs: a tab of General Notes that includes an input box for noting down the overall insights, and a tab of Evidence that shows the bookmarked charts and images. Detailed user comments can be appended to specific charts or images as fine-grained insights. Moreover, Vipera features LLM-powered auto-completion when users write in this view, further enhancing auditing efficiency. The prompts, bookmarked items, and existing notes will be injected into an LLM to generate the completion, which can be easily applied when users press the tab key on the keyboard.
@@ 4.2. User workflow and technical pipeline
Vipera has been developed and deployed as a web-based application. Figure 4 shows the workflow of Vipera [^2] All system prompts used in the pipeline are provided in the appendix.. After generating images from a user’s prompt, Vipera initiates the scene graph construction process, which is detailed in Figure 5. It starts by randomly sampling four images (or all, if fewer than four are available) to serve as a basis for the initial scene graph (Figure 5A). We chose this small sample size based on internal piloting, which indicated it struck an effective balance between capturing a diversity of concepts and ensuring low latency for the user, thus prioritizing usability and speed (DG1).
@@
For each sampled image, Vipera prompts an omni-modal LLM (Gemini 2.5 Flash in our implementation) to extract a tree-based scene graph that captures the image’s semantics (Figure 5B). The LLM is instructed to identify key physical objects and notable features and organize them into a tree structure. To provide a consistent high-level organization, we constrain the first level of the tree to “foreground” and “background” nodes in our study, which is an intuitive and widely applicable distinction. The resulting scene graphs from the four images are then merged, with their leaf nodes combined (Figure 5C). To avoid overwhelming the user (DG1), we ensure the scene graph’s size is approximately seven nodes as informed by the Miller’s Law. Thus, a maximum of five [^3] Our informal sensitivity checks suggested that minor variations in the scene graph parameters, such as sampling 3 or 6 images, or adjusting the node limit, did not qualitatively alter the conceptual breadth of the graph. leaf nodes are randomly selected for the final aggregated scene graph (Figure 5D), ensuring fairness and giving rarer but potentially insightful concepts an equal chance to appear in the initial scaffold. Note that the initial graph is not an exhaustive summary but rather a lightweight, interactive starting point. Subsequent user refinement during the auditing process inherently mitigates the impact of the initial random sampling.
@@
With the scene graph serving as a visual summary, users can add object or attribute nodes, where attribute nodes will serve as auditing criteria. When an attribute node is added, Vipera begins to evaluate and label the images using this criterion (Figure 4C). First, it updates the scene graph with the new node and ensures that all images within the new node’s scope are also included in the scope of its ancestor nodes up to the root. Then, it extracts a partial graph schema, which is a subset of the scene graph containing the new node and its path to the root. This schema is used to prompt an omni-modal LLM (GPT-5-mini in our implementation) to generate a label for each image within the specified scope. If the user has defined candidate values for the attribute, the LLM is constrained to choose from that list. The resulting labels are then aggregated, and their distribution is visualized as a stacked bar chart (Figure 4D) embedded into the added node using the D3.js library [^4] https://d3js.org/ (Figure 4E). Furthermore, users can bookmark the images and the charts within scene graph nodes at any time during auditing and comment about them in the note view for an auditing report (Figure 4F).
@@
Figure 5. The detailed technical pipeline for generating the initial scene graph. (A) A random subset is sampled from the images. (B) For each image, an individual scene graph is generated through an omni-modal LLM. (C) The result scene graphs are merged into one. (D) The result is pruned by keeping a maximum of five leaf nodes.
Same as the caption of this figure.
@@
Additionally, users can benefit from diverse sources of inspiration for iterative exploration of prompts and auditing criteria in Vipera, as illustrated by the blue arrows in Figure 4. Specifically, they may inspect the images carefully to get insights and new inspirations for prompts and criteria (Figure 4a). Such inspirations may also come from the scene graph, where the user either thinks with its structure about potential nodes to be added or examine the embedded charts for insights (Figure 4b). Finally, they may adopt the LLM-powered suggestions at any time for new auditing ideas (Figure 4c). Such an inspiration loop is built upon the artifacts from the data loop and serves as the engine for the user workflow, promoting iterative, mixed-initative, and comprehensive auditing.
@@ 4.3. Usage Scenario
To begin his audit, Bob enters the prompt “A cinematic photo of a doctor” into the input view and generates 15 images (Figure 2, step 1). Vipera populates the analysis view with the prompt and images, while displaying a corresponding scene graph that characterizes the semantics of the images (Figure 2, step 2). The graph organizes the content into foreground (e.g., doctor, nurse) and background (e.g., office, medical equipment) nodes.
@@
Afterwards, Bob felt too tired to come up with new auditing ideas and decided to refer to the AI suggestions. He first examined the Audit Analysis Support, which highlighted two images with a key difference and suggested adding a stethoscope attribute to the doctor node. Bob found the suggestion insightful and applied the suggestion (Figure 2, step 5). The resulting chart showed that some doctors were not wearing a stethoscope. To verify this, he hovered over the corresponding bar, which highlighted the specific images in the image view (Figure 3), allowing for rapid, coordinated sensemaking.
Next, Bob noticed the Prompt Suggestion recommending he replace “doctor” with “nurse” to explore a related profession. He accepted the prompt, and Vipera generated a new set of images (Figure 2, step 6). Crucially, the system also added a new nurse branch to the scene graph and automatically duplicated the existing criteria (gender and stethoscope) under it. With Vipera’s color-coded bar charts, Bob compared the results of these two prompts on the scene graph easily and acquired new insights into occupational stereotypes. From this point, Bob can continue his audit through multiple iterations of structured exploration and comparison.